# The Affective Bridge: Unifying Feature Representations for Speech Deepfake Detection

### v1

@@ III-B Results
This demonstrates that incorporating the emotion as a bridge strategy consistently improves the model’s discriminative capability compared to the pre-trained baseline across most datasets and feature types. Specifically, the EmoBridge configuration generally achieves lower EER and higher accuracy, indicating a trend particularly pronounced for systems using all features, where emotion-guided fusion effectively enhances the separability between genuine and fake speech. These results confirm that emotion-related feature fusion provides complementary information that benefits the deep representation for the deepfake detection task. Compared with the Emotion-pretrained HuBERT model, the performance before EmoBridge training is not dominant. We attribute this to the fact that the model was trained solely on the IEMOCAP dataset, which provides a limited and domain-specific representation of emotion. After fusing more diverse emotion knowledge through the EmoBridge strategy, the performance improves noticeably, suggesting that a broader and more balanced emotion distribution is crucial for model performance.
Moreover, across all datasets, the DL-raw features consistently benefit from the emotion as a bridge strategy, indicating that raw deep learning representations capture key cues for deepfake detection and that emotion information serves as an important complementary signal. The FoR corpus benefits the most from our strategy, showing substantial reductions in EER (e. g., from 0.082 to 0.045 for ASR features) along with notable accuracy gains. Improvements on the ITW dataset are also evident, particularly for deep-learned features, where the EER decreases from 0.033 to 0.020 and the accuracy increases from 0.969 to 0.982. These findings suggest that emotion-based bridging provides the greatest advantage in diverse or emotionally expressive conditions.
In contrast, the ASVSpoof shows comparable performance. To identify the subset where our strategy is most vulnerable, we further analyze model performance across different sources in the test set, shown in Table III.
TABLE III: Accuracy comparison across different sources and feature settings before and after the emotion as a bridge strategy. In this table, only the most noteworthy results are emphasised in bold.
Source ASR SV DL-raw Emotion
@@
A07 1.000 0.999 0.999 1.000 0.998 0.999 1.000 0.998
A08 0.995 1.000 0.999 0.999 0.949 0.995 0.988 0.999
A09 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
A10 0.999 0.996 0.995 0.996 0.842 0.857 0.998 0.977
A11 1.000 1.000 1.000 1.000 0.979 0.937 1.000 0.999
A12 1.000 0.998 1.000 1.000 0.999 0.991 1.000 0.998
A13 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
A14 1.000 1.000 1.000 1.000 0.987 0.990 1.000 1.000
A15 1.000 0.986 0.998 0.998 0.996 0.998 1.000 0.999
A16 0.999 0.995 0.997 0.996 0.994 0.994 0.999 0.993
A17 0.898 0.980 0.968 0.956 0.999 0.999 0.954 0.984
A18 0.956 0.996 0.994 0.990 0.948 0.969 0.793 0.975
A19 0.788 0.920 0.780 0.890 0.966 0.941 0.813 0.903
Real 0.695 0.520 0.749 0.486 0.860 0.884 0.820 0.564
The results indicate that the observed degradation mainly occurs for real audio under our strategy, likely due to the imbalanced training distribution affecting the emotion-fused features. From the DL-raw results, performance on the real subset remains higher, and overall accuracy improves, suggesting that emotion-enhanced features still contribute positively. Notably, audio generated by methods A17, A18, and A19 shows more consistent improvements than other source types, as these fakes are based on real human speech, whereas other methods rely on text or TTS synthesis. This suggests that human speech provides richer prosodic and emotional cues, resulting in more realistic fakes and more detectable emotional inconsistencies, while text- or TTS-based synthesis lacks natural emotional variation, yielding smaller gains from emotion-based modeling. Overall, our strategy remains effective, and as deepfake generation advances toward more expressive synthesis, fusing emotion-related information is expected to become very important for future detection systems.
III-C Explainability for model improvement
@@
Our approach yields performance gains, so we further investigate the model’s internal mechanism and feature representation. Specifically, we examine the mean attention values across time steps for different layers before and after applying our strategy, using the DL-raw features as an example, shown in Figure 2. The results show that the model after applying our strategy retains partial overlap with the original attention distribution, indicating that it preserves useful prior information, while also developing new attention regions where more affective cues are likely captured.
@@
Additionally, the previous claim is further supported through feature visualization using selected samples from the EmoFake dataset [zhao2024emofake]. These samples are controlled to share the same speaker, content, or emotion, allowing factor-wise comparison of feature representations. We analyze SV and ASR features, as shown in Figure 3, using t-SNE [maaten2008visualizing] for visualization.
After applying EmoBridge, the features cluster more distinctly by emotion, demonstrating the effectiveness of emotion feature fusion. Meanwhile, the original characteristics are largely preserved; for instance, in the SV task, samples with the same Roman numerals (indicating the same speaker) remain closer together, though the clusters are slightly less compact than before EmoBridge training. This suggests that our strategy avoids catastrophic forgetting during emotion fusion and prevents overfitting to specific tasks, thereby enhancing scalability.
The results show that the trends for emotion and deepfake detection representations are quite similar, indicating the importance of emotion features in deepfake detection. Lower layers capture more basic features, which are less informative for application-level tasks such as emotion recognition, while higher layers provide better representations for both emotion and deepfake detection.
IV Conclusion
We proposed the Emobridge strategy, a unified approach, to incorporate emotion information into pre-trained features. This approach not only preserves the original feature representations while adding affective cues, but also improves deepfake detection performance on multiple benchmarks. We consider emotion to be a more interpretable feature, perceivable by humans, which can serve as a bridge to obtain a unified feature representation. In future work, a unified representation for deepfake detection could be achieved by exploring methods to effectively combine representations from different pre-trained models through emotion.

### v2

@@ 3.2 Results
This demonstrates that incorporating the EmoBridge strategy consistently improves the model’s discriminative capability for deepfake compared to the pre-trained baseline across most datasets and feature types. Specifically, the EmoBridge configuration generally achieves lower EER and higher accuracy. Compared with the Emotion-pretrained HuBERT model, the performance before EmoBridge training is not dominant. We attribute this to the fact that HuBERT is trained on IEMOCAP, which provides a limited coverage of emotional expressions. By incorporating more diverse emotion knowledge, the detection performance improves noticeably, indicating that a broader and balanced emotion distribution is necessary.
Moreover, across all datasets, the DL-raw features consistently benefit from the EmoBridge, indicating that raw deep learning representations capture key cues for DFD and that guidance provided by emotion information is effective. The FoR corpus benefits the most from our strategy, showing substantial reductions in EER (e. g., from 0.082 to 0.045 for ASR features) along with notable accuracy gains. Improvements on the ITW dataset are also observed. Although the relative gains are modest, the large scale of the dataset means that these gains translate into a considerable number of additional correctly classified samples. These findings suggest that emotion-based bridging provides the greatest advantage in diverse or emotionally expressive conditions.
In contrast, comparable performance is observed on ASVSpoof. To identify where our approach is most vulnerable, we further analyse model performance across different sources within the test set. We also compare against two existing emotion-based methods: using the static emotion feature [conti2022deepfake] (results reported directly from the original paper), and a concatenating emotion feature with others (adapted from [mittal2020emotions]), which we combine HuBERT embeddings with DL-raw features (using our same experimental configuration of the models in Section 3.1). The results are shown in Table 3.
Table 3: Accuracy comparison across different sources and feature settings before and after the emotion as a bridge strategy. Column Static is from work [mittal2020emotions] (— means not provided or N/A) while Column Emotion is from our strategy using Hubert as original feature. In this table, only the most noteworthy results are emphasised in bold.
Source ASR SV DL-raw Emotion Static Concatenation
@@
A07 1.000 0.999 0.999 1.000 0.998 0.999 1.000 0.998 0.948 0.994
A08 0.995 1.000 0.999 0.999 0.949 0.995 0.988 0.999 0.988 0.962
A09 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
A10 0.999 0.996 0.995 0.996 0.842 0.857 0.998 0.977 0.900 0.965
A11 1.000 1.000 1.000 1.000 0.979 0.937 1.000 0.999 0.895 0.997
A12 1.000 0.998 1.000 1.000 0.999 0.991 1.000 0.998 0.890 1.000
A13 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 0.831 1.000
A14 1.000 1.000 1.000 1.000 0.987 0.990 1.000 1.000 0.763 0.952
A15 1.000 0.986 0.998 0.998 0.996 0.998 1.000 0.999 0.927 0.954
A16 0.999 0.995 0.997 0.996 0.994 0.994 0.999 0.993 0.898 0.988
A17 0.898 0.980 0.968 0.956 0.999 0.999 0.954 0.984 0.988
A18 0.956 0.996 0.994 0.990 0.948 0.969 0.793 0.975 0.869
A19 0.788 0.920 0.780 0.890 0.966 0.941 0.813 0.903 0.905
Real 0.695 0.520 0.749 0.486 0.860 0.884 0.820 0.564 0.832
The results indicate that the relatively smaller gains observed on ASVSpoof2019 LA can be attributed to the highly imbalanced distribution of bona fide and spoofed samples in this dataset, where real speech constitutes only a small portion of the training data. Nevertheless, the DL-raw results still achieve higher performance on the real subset and improved overall accuracy, indicating that emotion-enhanced representations remain beneficial. Audio generated by methods A17–A19 shows more consistent improvements than other source types, likely because these methods are from real human speech, which contains richer prosodic and emotional variation than TTS-based synthesis. This observation does not suggest a limitation of the proposed framework, but highlights the importance of bona fide diversity when emotion is used as an alignment signal.
3.3 Effects of DFD classifier
We select SVM as the main classifier for simplicity, although other classifiers could also be used. Previous work shows that the relative performance of different feature representations remains consistent across classifier choices [chakravarty2024lightweight], even though absolute scores may vary. To further validate the robustness of our approach, we conduct an ablation study on the FoR dataset using a two-layer MLP classifier (hidden dimension 512 with a 2-dimensional output), as shown in Table 4. The results show similar performance across feature types, demonstrating the robustness of our method.
Table 4: EER and accuracy between the Pre-trained model and the model with the EmoBridge strategy on FoR with MLP as classifier.
Dataset Model Pre-trained Emotion fused
EER $\downarrow$ Acc $\uparrow$ EER $\downarrow$ Acc $\uparrow$
FoR openSMILE .385 .554 .265 .645
Emotion .143 .866 .083 .915
SV .146 .864 .087 .910
ASR .088 .898 .066 .952
DL-raw .095 .899 .077 .943
3.4 Preservation of speech representation
Additionally, we evaluated the preservation of speech representation via feature visualization using selected samples from the EmoFake [zhao2024emofake]. These samples are controlled to share the same speaker, content, or emotion, allowing factor-wise comparison of feature representations. We analyze SV and ASR features, as shown in Figure 2, using t-SNE [maaten2008visualizing] for visualization.
After applying EmoBridge, the features cluster more distinctly by emotion, demonstrating the effectiveness of emotion guidance. Meanwhile, the original characteristics are largely preserved; for instance, in the SV task, samples with the same Roman numerals (indicating the same speaker) remain closer together, although the clusters are slightly less compact than before EmoBridge training. Directly evaluating ASR performance via word error rate after EmoBridge is non-trivial, as ASR systems rely on a tightly coupled encoder–decoder architecture, and re-training or re-adapting the decoder would introduce additional confounding factors. We therefore assess representation preservation through encoder-level, task-agnostic evaluation. Specifically, we compare speaker verification performance in a zero-shot setting. Using the same settings as in [aldeneh2024remove], the models before and after EmoBridge achieve EERs of 0.313 and 0.325, respectively, on VoxCeleb [nagrani2020voxceleb], indicating no huge degradation. This suggests that the proposed strategy avoids fully altering the original semantic information and prevents overfiting to a specific downstream task, enhancing scalability.
3.5 Explainability for model improvement
@@
Our approach yields performance gains, so we further investigate the model’s internal mechanism. Specifically, we examine the mean attention values across time steps for different layers before and after applying our strategy, using the DL-raw features as an example, shown in Figure 3. The results show that the model after applying our strategy retains partial overlap with the original attention distribution, indicating that it preserves useful prior information, while also developing new attention regions where more affective cues are likely captured.
@@
4 Conclusion
We propose EmoBridge, a unified strategy that leverages emotion information as a constraint to guide pre-trained feature embeddings. This approach preserves the original feature representations while injecting affective cues, enabling the model to better expose discriminative patterns for DFD and consistently outperform baseline methods across multiple benchmarks. Our experiments show that emotion information provides the effective guidance for improving representation quality in this framework. Future work could explore integrating multimodal features with emotion-aware guidance for DFD.
5 Generative AI Use Disclosure
We only use Generative AI to check the grammar issue of our paper.